Consider linear operator T:C^3-->C^3 def by T(z1,z2,z3)=(z1+iz2, iz1-2z2, -iz2+z3). Check T* and check if T is self adjoined. check if T is unitary
1
Expert's answer
2020-03-23T10:14:56-0400

Here matrix representation of T with respect to standard basis "\{ (1,0,0), (0,1,0), (0,0,1)\}" over the field "C" is "A="

"\begin{pmatrix}
 1& i& 0 \\ 
i & -2&0\\
0&-i&1
\end{pmatrix}"

Now , "T^*=" Transpose conjugate of matrix representation of "T"

"=A^H"

"=" "\begin{pmatrix}
 1&-i&0 \\
 -i&-2&0\\
0&i&1

\end{pmatrix}"

Since ,"A\neq A^H"

Hence, "T" is not a self adjoint operator.

To check unitary operators ,

Consider "A.A^H=" "\begin{pmatrix}
 2&-3i&-1\\
3i&5&-2i\\
-1&2i&2
\end{pmatrix}" "\neq I"

Where "I" stands for identity matrix.

Therefore,"T" is not a unitary operator.